Air is pumped into a spherical balloon at the rate of 100 cm3/s. How fast is the diameter increasing when the diameter is 10 cm? 


1
Expert's answer
2022-05-16T17:01:11-0400

"V=4/3\pi r^3"

"r=0.5 D"

"V=4/3\pi (0.5D)^3=1/6 \pi D^3"

"dV/dt=1/2\pi D^2 dD/dt"

"dD/dt=(dV/dt)/(0.5\pi D^2)=100/(0.5(10)3.14)=6.37 cm/s"